OpenStudy (anonymous):

Why does the moderator name Khan keep deleting my question, when i ask for HELP not for the answers?

OpenStudy (anonymous):

idk.

OpenStudy (anonymous):

odds are, he/she will delete this one too. and the mods think they can delete what they want unfortunatly. they have done the same to me.

OpenStudy (anonymous):

i wish they can stop, and go find a life instead of doing something they call a "JOB" not all mobs lol some of them are really nice.

OpenStudy (anonymous):

^*mods

OpenStudy (anonymous):

Mods

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

lol yes, thank you for the lesson lol :)

OpenStudy (anonymous):

:) so what qustion was it ? maybe i can help

OpenStudy (anonymous):

here it go 1. d2 + 18d + 81 (1 point) (d + 9)(d – 9) (d + 9)2 (d – 81)(d – 1) (d – 9)2 2. d2 – 22d + 121 (1 point) (d + 11)2 (d – 11)2 (d – 11)(d + 11) (d –121)(d –1) 3. r2 – 49 (1 point) (r – 7)(r + 7) (r + 7)(r + 7) (r – 7)(r – 7) (r – 7)(r + 9) 4. k2 – 81h2 (1 point) (k – 9h2)(k + 9) (k + 9h)(k + 9h) h2(k + 9)(k – 9) (k + 9h)(k – 9h) 5. 25x2 + 40x + 16 (1 point) (5x + 4)2 (5x – 4)2 (5x + 4)(5x – 4) (25x + 4)(x + 4)

OpenStudy (anonymous):

well u should definitly space it out more. its a lil hard to read. try posting one question at a time. and im not good at algebra at all but i can find someone who is :)

OpenStudy (anonymous):

1. d2 + 18d + 81 (1 point) (d + 9)(d – 9) (d + 9)2 (d – 81)(d – 1) (d – 9)2

OpenStudy (anonymous):

i meant when u go to post a question, post one at a time. and i found this answer searching around (d+9)^2

OpenStudy (anonymous):

(d+11)^2 for the next one (again im jus giving answers im finding while searching around)

OpenStudy (anonymous):

(r+7)(r-7) this one for the third

OpenStudy (anonymous):

for the fourth k^2-81h^2 =(k)^2-(9h)^2 =(k+9h)(k-9h)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

the fifth 25x^2 + 40x + 16 = (5x + 4)^2

OpenStudy (anonymous):

again, im jus searchinga round and giving u the answers i find. i could be wrong.

OpenStudy (anonymous):

no u wasnt

OpenStudy (rosedewittbukater):

Most moderators are just teachers who want to prevent kids from actually understanding the subject because they believe they are teaching us well when they cram in 2 weeks of learning into an hour session that just end up confuse us more.

OpenStudy (anonymous):

1) (d + 9)2

OpenStudy (anonymous):

2) (d + 11)2

OpenStudy (anonymous):

3) r – 7)(r + 7)

OpenStudy (anonymous):

4) (k + 9h)(k – 9h)

OpenStudy (anonymous):

5) (5x + 4)2